OpenStudy (anonymous):

4(-x^2-25)/(x^2-25)^2=0 solve for x. Help!

OpenStudy (anonymous):

Well you need where the numerator is 0. Therefore we have: \[-4 \frac{x^2+25}{(x^2-25)^2}=0 \implies \frac{(x+5i)(x-5i)}{(x+5)^2(x-5)^2} = 0 \implies x= \pm 5 i\] where i is the principal square root of -1, \[i = \sqrt{-1}\]